OpenStudy (anonymous):

What is the static deflection of a spring having stiffness of 2000 N/m under the load of an object with mass of 10 kg? Recall that the Earth’s gravitational acceleration is approximately g=9.8m/sec2 and F=m×g.

OpenStudy (whpalmer4):

F = -kx counterbalances F = mg 2000 N/m * x = 10 kg*9.8 m/s^2 = 98 N x = 98 N / 2000 N/m = 0.049 m